# Practical indistinguishability in a gene regulatory network inference problem, a case study

### v1

@@ 1 Abstract
Computationally inferring mechanistic insights and underlying control structures from typical biological data is a challenging pursuit. The technical reasons for this are multifaceted–and we delve into them in depth here, but they are easy to understand and involve both the data and model development. Even the highest-quality experimental data come with challenges. There are always sources of noise, a limit to how often we can measure the system, and we can rarely measure all the relevant states that participate in the full underlying complexity. There are usually sources of uncertainty in model development, which give rise to multiple competing model structures. To underscore the need for further analysis of structural uncertainty in modeling, we use a meta-analysis across six journals covering mathematical biology and show that a huge number of mathematical models for biological systems are developed each year, but model selection and comparison across model structures appear to be less common. We walk through a case study involving inference of regulatory network structure involved in a developmental decision in the nematode, Pristonchus pacificus. We first examine the practical indistinguishability of a model structure, or the ability to uniquely infer the structure given the data, across a wide range of synthetic data regimes by refitting both the true model structure and several misspecified models. We then use real biological data and compare across 13,824 models–each corresponding to a different regulatory network structure, to determine which regulatory features are supported by the data across three experimental conditions. We find that the best-fitting models for each experimental condition share a combination of features and identify a regulatory network that is common across the model sets for each condition. This model is capable of describing the data across the experimental conditions we considered and exhibits a high degree of positive regulation and interconnectivity between the key regulators, eud-1, sult-1, and nhr-40. While the biological results are specific to the molecular biology of development in Pristonchus pacificus, the general modeling framework and underlying challenges we faced doing this analysis are widespread across biology, chemistry, physics, and many other scientific disciplines.
@@ 2 Introduction
Data-driven modeling, or the act of developing and parameterizing mathematical models from experimental data, can be a powerful tool to explain, understand, and make predictions about how the underlying system operates. However, parameterizing mathematical models through numerical optimization is a challenging endeavor in its own right, due to properties of biological data and structural uncertainty. Biological experiments across fields often result in data that include sources of intrinsic and extrinsic noise, are sampled infrequently in time, and usually, not all relevant components of the underlying system can be measured directly, even with the most cutting-edge experimental equipment and designs. As a result, we might have well-founded descriptions of some components of the biological system in a model, but multiple possible model descriptions for the less resolved or well-understood processes. This leads to multiple competing structures, and is known as structural uncertainty. For example, it may be known that a single gene embedded within a complex network is regulated by a transcription factor, but unclear if the regulation is positive or negative in nature. Structural uncertainty is widespread in biological modeling, and subtle changes in the model structure can influence the results. This topic is rarely explored in depth due to the computational burden of attempting to distinguish between model structures in a data-driven manner. Distinguishing between multiple model structures requires navigating high-dimensional cost function landscapes–which can be challenging for even a single model, to determine the best fit of each model structure to the data through numerical optimization.
@@
Fitting even a single model to data through numerical optimization can be challenging for several reasons, including generalized features of the cost function landscape associated with models of biological systems and both structural and practical identifiability issues, which can emerge based on properties of typical biological data. Broadly, cost function landscapes for biological systems are known to feature long, flat canyons, where parameter values can widely vary but the model behavior stays qualitatively the same [1]. These landscapes are hard to fully numerically explore, even when using synthetic data that is noise-free. In the context of realistic experimental data, the cost function landscapes can easily become rippled or distorted, making it more difficult to perform parameter estimation. The properties of experimental data can also impact both structural and practical identifiability of the model. Typically, not all relevant states of the system can be measured, leading to hidden state variables. Hidden states can lead to structural identifiability issues, which occur when the the parameters of the model cannot be uniquely inferred based on the structure of the model alone. Both low temporal data sampling and sources of noise can lead to practical identifiability issues, which occur when the parameters of the model cannot be uniquely determined given the data. Both structural and practical indentifiability issues lead to a lack of uniqueness in the estimated parameters, which can impact the ability to use the model for explanation or prediction. Further complications emerge when we consider multiple model structures to take into account structural uncertainty.
@@
Structural uncertainty leads to multiple competing models, and every change to the structure of the model impacts the geometry of the cost function landscape, potentially altering the ease of parameter estimation. Adding additional parameters to a model increases the dimensionality of the cost function landscape, whereas imposing changes to the functional form of the model that keep the number of parameters fixed, for example, the difference between a Hill Function that describes activation or repression of a gene transcription, can distort or bend the geometry of the cost function landscape, so it is important to consider the sensitivity of the modeling results as a function of all possible model structures that emanate from structural uncertainty, especially in our current scientific moment, when so many mathematical models of biological systems are being produced.
@@
A huge number of mathematical models are being developed and published each year (see Figure 1). To get a crude and imperfect estimate of the lower bound for how many articles are comparing across model structures, we went to each of the journals’ websites that we considered and searched for “Akaike Information Criterion,” to give a lower bound (see Methods). Akaike Information Criterion is a classical and widely used statistical approach to model selection [2]. Across all time, we found between 40 and 133 results for this search in more classical mathematical biology journals  [^1] We considered Bulletin of Mathematical Biology, the Biophysical Journal, the Journal of Biological Rhythms, the Journal of Mathematical Biology, and the Journal of Theoretical Biology to be more focused on classical mathematical biology., but 2,649 results (out of 11,461 total results, as indexed on PubMed) in PLOS Computational Biology, which has more of a focus on data-driven modeling. It seems reasonable to conclude that at least 23% of studies in PLOS Computational Biology contain some degree of model selection, but comprehensive model structure comparison is far from common (see Methods for further details). A handful of recent studies have begun to carefully consider the influence of changes in model structure on modeling results, as we explore here. For example, Bayesian model selection was used to compare across 4,208 possible models of interactions between microbial communities [3]. Model selection on more than a handful of models is not yet common across mathematical biology, and more efforts to examine sources of structural uncertainty are important. Here, we take a maximalist approach and grapple with a biological problem that has a high degree of structural uncertainty. We fit a large family of 13,824 Ordinary Differential Equation (ODE) models, each model corresponding to a distinct regulatory network structure, to an existing gene expression data set to infer the currently unknown structural elements of a regulatory network that underpins a developmental decision in the nematode Pristionchus pacificus.
@@
Two genes in the network have been identified as “switch” genes. When either gene is mutated, worms exhibit a 100% penetrant mouth-form phenotype regardless of environmental conditions, and when these genes are overexpressed, worms exhibit the opposite phenotype [11, 13, 14]. These two genes code for enzymes with related, but opposite functions: (eud-1) encodes a sulfatase, whereas (seud-1/sult-1, hereafter referred to as sult-1), encodes a sulfotransferase. Sulfotransferases and sulfatases are enzymes with opposite biochemical functions; Sulfotransferases add and sulfatases remove sulfate groups from molecular substrates within the cell. Sulfation affects the activity of signaling molecules, is required for the metabolism and lysosomal degradation of many compounds, and generates diverse extracellular ligands [15, 16]. eud-1 and sult-1 expression have opposite effects on mouth-form development, indicating that the relative abundance of these enzymes and their unidentified substrates determines the mouth-form phenotype. Furthermore, a recent RNA-seq experiment revealed that eud-1 and sult-1 are the only genes in the network whose expression is dependent on the environment during the critical window for mouth-form development [17]. Thus, understanding the regulatory connections of these two genes is critical to understanding the logic of mouth-form plasticity. As an experimentally tractable system where the components of the gene regulatory network are known, determining their connectivity can inform plasticity writ large.
@@
Suppressor screens have identified several genes that operate downstream of eud-1 and sult-1, although the structure and regulatory connections between them are also not well understood. The nuclear hormone receptor nhr-40 is particularly interesting because it is a transcription factor, and it yields opposite phenotypes depending on the mutation (i.e., loss-of-function and gain-of-function alleles). To keep our analysis computationally tractable on a high-performance computational cluster, we explore a large family of effective regulatory network models involving eud-1, sult-1, and nhr-40, which are thought to be the key players of the regulatory network underpinning the developmental decision in the nematode P. pacificus. Beyond the developmental biology questions we pursue here, our general approach to mechanistic inference from typical experimental data can be applied to modeling contexts across biology where structural uncertainty is an issue.
@@
Figure 2: A: The regulatory network that underpins the dimorphic developmental mouthform decision in P. pacificus is thought to involve eud-1, nhr-40, sult-1, but the regulatory structure of the network is unknown. We use gene expression data sampled in three different experiments to identify key features of the regulatory network. B: We fit a family of ODE models to the gene expression data from the experiments. C: We assess the model fits to data across model structures. D: We then infer what features of the model structure lead to acceptable fits to the data, which we refer to as model sets.
@@
We fit 13,824 distinct ODE model structures—each corresponding to a distinct effective regulatory network involving eud-1, sult-1 and nhr-40, to RNA-seq data from three experimental conditions (see Sections 6.2.1, 6.3, 6.4, 6.5). Using our maximalist computational approach, we resolve some features of the effective regulatory network, but we cannot completely distinguish between models given the data. We identify model sets, or combinations of features that are shared across the models generating acceptable fits to each dataset (see Section 6.6). Similarly, we can rule out a host of possible regulatory network structures. We find a highly interconnected regulatory network structure dominated by positive regulation that can explain the data across each of the three experimental conditions, suggesting a possible mechanism that underpins the mouth-form decision in Pristionchus pacificus. More work is needed to experimentally verify the proposed mechanism and, more broadly, refine our model sets. The results we find here could be used to guide experiments aimed at achieving a mechanistic understanding of regulation during developmental decisions. This study examines practical indistinguishability, a generalized form of sensitivity analysis across model structure given a data set, which is rarely considered in modeling studies but is increasingly important given the growing number of models being created. Our analysis allows us to probe the limits of what we can mechanistically infer, and is broadly applicable across scientific disciplines, including biology, chemistry, and physics.
@@ 3 Results
To ensure that the model structures we considered were not a source of non-unique parameter fits, we tested representatives from our large family of models for structural identifiability using STRIKE-GOLDD, and found they were not structurally identifiable, due to the presence of a scaling symmetry [18]. Scaling symmetries are a type of algebraic transformation of parameters and variables of a model that leave the model invariant, meaning it produces the same dynamic behavior in the observables. Consider the following example model from our family of regulatory network models with the scaling symmetry highlighted in blue. From visual inspection, the scaling symmetry cancels from each equation of the model, meaning any value of $\xi$ produces the same mRNA dynamics.
@@
[EUD-1 protein], Unobs.: ${\color[rgb]{0,0,1}\xi}\dfrac{dE}{dt}={\color[rgb]{0,0,1}\xi}k_{t}M_{e}-{\color[rgb]{0,0,1}\xi}k_{-p}\hat{E}$, $\hat{E}(0)=\hat{E_{o}}$ (2)
@@
[NHR-40 protein], Unobs: ${\color[rgb]{0,0,1}\xi}\dfrac{dN}{dt}={\color[rgb]{0,0,1}\xi}k_{t}M_{n}-{\color[rgb]{0,0,1}\xi}k_{-p}N,$ $N(0)=\hat{N_{o}}$ (4)
@@
[NHR-40 protein], Unobs: ${\color[rgb]{0,0,1}\xi}\dfrac{dN}{dt}={\color[rgb]{0,0,1}\xi}k_{t}M_{n}-{\color[rgb]{0,0,1}\xi}k_{-p}N,$ $N(0)=\hat{N_{o}}$
@@
To eliminate the structural identifiability issue and break the scaling symmetry, we set $\xi=\dfrac{1}{k_{t}}$, which rescales the protein states and the half-max parameters by the translation rate. This is reasonable, as we know that the translation rate must be non-zero and has the benefit of reducing the dimension of the parameter space by one. We then checked a subset of the reparameterized models to ensure the resulting models were structurally identifiable, using an alternative algorithm [19]. For example, Eqs. (7)-(12) are a structurally identifiable version of Eqs. (1)-(6). We corrected all 13,824 models for this scaling symmetry ahead of parameter estimation (see Sections 6.3 and 6.4). Although we cannot verify the structural identifiability of all remaining models due to computational limitations, removing at least this ubiquitous symmetry should improve the reliability of parameter estimation.
@@
Before fitting our family of models to the experimental data, we performed a series of synthetic tests using Eqs. (7)-(12) to understand the limits of inference in different data regimes. As Eqs. (7)-(12) are structurally identifiable, any parameter identifiability issues should be practical in nature, meaning they emerge from the data quality rather than from symmetries in the structure of the equations. To do this, we sampled synthetic data from known parameters and varied the sampling frequency and noise level. We then refit the true model structure to the synthetic data to get a sense of how well we can recover the ground truth parameters. We also fit several “misspecified model structures,” or models with structural differences from the true model structure, to the synthetic data to understand the practically indistinguishable of Model 11574 from other models across different data regimes. We refer to practical indistinguishability as the ability to uniquely infer the structure given the data. Practically indistinguishable was less examined in our search of the literature, perhaps due to a combination of its computational nature or the challenge of interpreting the results across multiple, competing model structures that emerge from structural uncertainty, though we did find some examples, such as [3].
@@
To determine how well we can recover ground truth parameters, we simulate synthetic data across a range of sampling rates corresponding to a few observations across the course of development to hourly sampling and a broad range of noise levels from Model 11574 (of 13,824) and attempt to fit the true model structure to the data (Figure 3, B top). We select this model because it produces dynamics similar to those observed in the data, which we discuss in subsequent sections. Model 11574 has the form
[eud-1 mRNA], Observed: $\dfrac{dM_{e}}{dt}=k_{t_{e_{1}}}\left(\dfrac{\hat{N}}{\hat{K_{1}}+\hat{N}}\right)\left(\dfrac{\hat{S}}{\hat{K_{4}}+\hat{S}}\right)-k_{-m}M_{e},$ $M_{e}(0)=M_{eo}$ (7)
[scaled EUD-1 protein], Unobserved: $\dfrac{d\hat{E}}{dt}=M_{e}-k_{-p}\hat{E}$, $\hat{E}(0)=\hat{E_{o}}$ (8)
[nhr-40 mRNA], Observed: $\dfrac{dM_{n}}{dt}=k_{t_{n_{1}}}\left(\dfrac{\hat{N}}{\hat{K_{2}}+\hat{N}}\right)-k_{-m}M_{n},$ $M_{n}(0)=M_{no}$ (9)
[scaled NHR-40 protein], Unobserved: $\dfrac{d\hat{N}}{dt}=M_{n}-k_{-p}\hat{N}$, $\hat{N}(0)=\hat{N_{o}}$ (10)
[sult-1 mRNA], Observed: $\dfrac{dM_{s}}{dt}=k_{t_{s_{1}}}\left(\dfrac{\hat{E}}{\hat{K_{3}}+\hat{E}}\right)^{2}-k_{-m}M_{s},$ $M_{s}(0)=M_{so}$ (11)
[scaled SULT-1 protein], Unobserved: $\dfrac{d\hat{S}}{dt}=M_{s}-k_{-p}\hat{S}$, $\hat{S}(0)=\hat{S_{o}}$, (12)
@@
To analyze the impact of experimental noise on the practical distinguishability of this model from others, we needed to choose an error model. We observed that the noise was higher when the mean of the data was higher, potentially indicating the presence of signal-dependent noise, which is commonly observed in biology, for example, see [20]. Based on this, a multiplicative error model is the simplest reasonable model we can assume, as there was not enough data to empirically estimate the error distribution and there was no straightforward first-principles approach based on knowledge of the system. We assumed a data-centered Gaussian with standard deviation proportional to the mean and generate synthetic data for a wide range of sampling rates (row in Figure 3B) and noise levels (column in Figure 3B). More negative log-likelihood values (lighter colors) correspond to better fits (see Methods). Using this error model and the known sampling rate, we can approximate the experimental data regime (yellow box, Figure 3). When we fit the known model structure to synthetic data in the approximate experimental data regime, we get a reasonable fit to the data, though the exact parameter values are not recovered (Figure 3, B top, yellow box and C top). Based on these synthetic tests, we do not expect each individual model to have practically identifiable parameters.
@@
To understand if we can uniquely infer the structure of a given model from data sampled in the relevant experimental regime—or how practically distinguishable Model 11574 is from other model structures, we explore fitting misspecified models to the synthetic data. We select three alternative structures or misspecified models to fit to the synthetic data generated by Model 11574. We selected Models 7308, 353, and 1 of the 13K structures, as they vary in how closely related their structure is to Model 11574, the true model structure in this synthetic test (Figure 2 A). Model 7308 and Model 11574 share several regulatory features, whereas Models 353 and 1 have completely different regulatory structures from Model 11574 and vary in their complexity. Model 353 is much more complex than Model 1. In the relevant experimental data regime, we see that Model 7308 gives a reasonable fit to synthetic data generated from Model 11574 and has relatively tightly constrained uncertainty (Figure 2 C, second row), whereas Models 353 and 1 give a much poorer and more uncertain fit (Figure 2 C, third and fourth rows). In particular, the strongly mispecified models poorly describe the dynamics of nhr-40 mRNA over the entire course of development, as the uncertainty covers the entire possible range of the data. While we do not expect the parameters to be practically identifiable in the approximate experimental data regime, it is clear that some model structures are capable of generating a reasonable fit to the data, while others are not. iven the properties of the experimental data, we expect to be able to resolve some features of the network structure partially, but we do not expect this system to exhibit full practical distinguishability between all models.

### v2

@@ 1 Abstract
Determining mechanistic models of gene regulation, especially underlying phenotypic variation, is a central goal of both mathematical biology and modern evolutionary biology. Progress has been made by analyzing high-throughput sequencing data, yet several challenges, involving both common characteristics of experimental data and the model development process, remain that limit the discovery of general principles. Even the highest-quality experimental data come with challenges. There are always sources of noise, a limit to how often we can measure the system in time, and it is impossible to measure all the relevant states that participate in the full underlying complexity. Additionally, there are usually sources of uncertainty in the underlying biological mechanisms, which give rise to multiple competing model structures. To underscore the need for further analysis of structural uncertainty in modeling, we use a meta-analysis across six journals covering mathematical biology and show that a huge number of mathematical models for biological systems are developed each year, but model selection and comparison across model structures appear to be less common. We then walk through a case study involving inference of regulatory network structure involved in a developmental decision in the nematode, Pristonchus pacificus, which exhibits an evolutionary novelty of mouth-dimorphism (predator vs. bacterivore). We first perform a careful study of the practical indistinguishability of a model structure, or the ability to uniquely infer the structure given the data, across a wide range of synthetic data regimes by refitting both the true model structure and several misspecified models. We then fit 13,824 distinct regulatory network models to gene expression data from three experimental conditions to determine which regulatory features are supported by the data. We discover model sets, or collections of models with shared regulatory network features that best fit the data, for each of the three experiments we considered and identify a regulatory network in the intersection of the three model sets. This model describes the data across the experimental conditions and exhibits a high degree of positive regulation and interconnectivity between the key regulators, eud-1, sult-1, and nhr-40. While the biological results are specific to the molecular biology of development in Pristonchus pacificus, the comparative modeling framework introduced here can be applied to other systems of gene regulation in an evolutionary developmental (evo-devo) context.
@@ 2 Introduction
Developing and parameterizing mathematical models from experimental data, can be a powerful tool to explain, understand, and make predictions about how the underlying system operates. However, parameterizing mathematical models through numerical optimization is a challenging endeavor in its own right, due to properties of biological data and structural uncertainty. Biological experiments across fields often result in data that include sources of intrinsic and extrinsic noise, are sampled infrequently in time, and usually, not all relevant components of the underlying system can be measured directly, even with the most cutting-edge experimental equipment and designs. As a result, we might have well-founded descriptions of some components of the biological system in a model, but multiple mechanistic hypotheses for the less resolved or well-understood components. This leads to multiple competing model structures, and is known as structural uncertainty. For example, it may be known that a single gene embedded within a complex network is regulated by a transcription factor, but unclear if the gene experiences activation or repression by the transcription factor. Structural uncertainty is widespread in biological modeling, and subtle changes in the model structure can influence the results. This topic is rarely explored in depth due to the computational burden of distinguishing between model structures in a data-driven manner. Distinguishing between multiple model structures requires navigating high-dimensional cost function landscapes for each model–a process that can be challenging for even a single model, to determine the best fit of each model structure to the data through numerical optimization.
@@
Fitting even a single model to data through numerical optimization can be challenging for several reasons, including global and location geometry of the cost function landscape associated with models of biological systems and both structural and practical identifiability issues, which can emerge based on properties of typical biological data. Broadly, cost function landscapes for biological systems are known to feature long, flat canyons, where parameter values can widely vary but the model behavior stays qualitatively the same [1]. These landscapes are hard to fully numerically explore, even when using synthetic data that is noise-free. In the context of realistic experimental data, the cost function landscapes can easily become rippled or distorted, making it more difficult to perform parameter estimation. The properties of experimental data can also impact both structural and practical identifiability of the model. Typically, not all relevant states of the system can be measured, leading to hidden state variables. Hidden states can lead to structural identifiability issues, which occur when the the parameters of the model cannot be uniquely inferred based on the structure of the model alone. Both low temporal data sampling and sources of noise can lead to practical identifiability issues, which occur when the parameters of the model cannot be uniquely determined given the data. Both structural and practical identifiability issues lead to a lack of uniqueness in the estimated parameters, which can impact the ability to use the model for explanation or prediction. See [2] for a recent review of identifiability analysis of linear and nonlinear models and the associated computational considerations. Further complications emerge when we consider multiple model structures to take into account structural uncertainty.
@@
Every change to the structure of the model impacts the geometry of the cost function landscape, potentially altering the ease of parameter estimation. Adding additional parameters to a model increases the dimensionality of the cost function landscape, whereas imposing changes to the functional form of the model that keep the number of parameters fixed, for example, the difference between a Hill Function that describes activation or repression of a gene transcription, can distort or bend the geometry of the cost function landscape, so it is important to consider the sensitivity of the modeling results as a function of all possible model structures that emanate from structural uncertainty, especially in our current scientific moment, when so many mathematical models of biological systems are being produced.
@@
A huge number of mathematical models of biological systems are being developed and published each year (see Figure 1). To get a crude and imperfect estimate of the lower bound for how many articles are comparing across model structures, we went to each of the journals’ websites that we considered and searched for “Akaike Information Criterion,” to give a lower bound (see Methods). Akaike Information Criterion is a classical and widely used statistical approach to model selection [3]. Across all time, we found between 40 and 133 results for this search in more classical mathematical biology journals  [^1] We considered Bulletin of Mathematical Biology, the Biophysical Journal, the Journal of Biological Rhythms, the Journal of Mathematical Biology, and the Journal of Theoretical Biology to be more focused on classical mathematical biology., but 2,649 results (out of 11,461 total results, as indexed on PubMed) in PLOS Computational Biology, which has more of a focus on data-driven modeling. It seems reasonable to conclude that at least 23% of studies in PLOS Computational Biology contain some degree of model selection, but comprehensive model structure comparison is far from common (see Methods for further details).
Some modeling studies of biological and ecological systems have attempted to carefully consider the influence of changes in model structure on modeling results, as we do here. A 2006 study fit a suite of seven differential equation models to Chronic Wasting Disease (CWD) outbreak data to examine different modes of transmission of Chronic Wasting Disease in deer populations [4], though there is some question as to whether all the model structures considered were structurally identifiable [5]. A similar study was performed by Lee and co-authors to examine the influence of five different model structures on epidemiological modeling and forecasting of Cholera transmission [6]. A more recent study used Bayesian model selection to compare across 4,208 possible models of interactions between microbial communities [7]. A recent article used recurrent neural networks to avoid a combinatorial search over all possible model structures for gene regulatory network inference, but the method appears to require higher temporal sampling than the data set we will examine here [8]. Additionally, a large array of model selection methods have been developed for data-driven model discovery of partially observed systems and avoid a comparison of all possible model structures, but typically these methods are demonstrated on higher temporal sampling [9, 10, 11, 12]. Here, we take a maximalist approach and grapple with a challenging regulatory network inference problem. We fit a large family of 13,824 Ordinary Differential Equation (ODE) models, each model corresponding to a distinct regulatory network structure, to an existing gene expression data set to infer the currently unknown structural elements of a regulatory network that underpins a developmental decision in the nematode Pristionchus pacificus.
@@
Two genes in the network have been identified as “switch” genes. When either gene is mutated, worms exhibit a 100% penetrant mouth-form phenotype regardless of environmental conditions, and when these genes are overexpressed, worms exhibit the opposite phenotype [20, 22, 23]. These two genes code for enzymes: (eud-1) encodes a sulfatase, whereas (seud-1/sult-1, hereafter referred to as sult-1), encodes a sulfotransferase. Sulfotransferases and sulfatases are enzymes with opposite biochemical functions; Sulfotransferases add and sulfatases remove sulfate groups from molecular substrates within the cell. Sulfation affects the activity of signaling molecules, is required for the metabolism and lysosomal degradation of many compounds, and generates diverse extracellular ligands [24, 25]. eud-1 and sult-1 expression have opposite effects on mouth-form development, indicating that the relative abundance of these enzymes and their unidentified substrates determines the mouth-form phenotype. Furthermore, a recent RNA-seq experiment revealed that eud-1 and sult-1 are the only genes in the network whose expression is dependent on the environment during the critical window for mouth-form development [26]. Thus, understanding the regulatory connections of these two genes is critical to understanding the logic of mouth-form plasticity. As an experimentally tractable system where the components of the gene regulatory network are known, determining their connectivity can inform plasticity writ large.
@@
Suppressor screens have identified several genes that operate downstream of eud-1 and sult-1, although the structure and regulatory connections between them are also not well understood. The nuclear hormone receptor nhr-40 is particularly interesting because it is a transcription factor, and it yields opposite phenotypes depending on the mutation (i.e., loss-of-function and gain-of-function alleles). To keep our analysis computationally tractable on a high-performance computational cluster, we explore a large family of effective regulatory network models involving eud-1, sult-1, and nhr-40, which are thought to be the key players of the regulatory network underpinning the developmental decision in the nematode P. pacificus. Beyond the developmental biology questions we pursue here, our general approach to mechanistic inference from typical experimental data can be applied to modeling contexts across biology where structural uncertainty is an issue and the properties of the experimental data makes algorithmic model selection infeasible.
@@
Figure 2: A: The regulatory network that underpins the dimorphic developmental mouthform decision in P. pacificus is thought to involve eud-1, nhr-40, sult-1, but the regulatory structure of the network is unknown. We use gene expression data sampled in three different experiments to identify key features of the regulatory network. B: We fit a family of ODE models to the gene expression data from the experiments. C: We assess the model fits to data across model structures. D: We then find the model set, or collections of models with shared regulatory network features that best fit the data.
@@
We fit 13,824 distinct ODE model structures—each corresponding to a distinct effective regulatory network involving eud-1, sult-1 and nhr-40, to RNA-seq data from three experimental conditions (see Figure 2 and Sections 6.2.1, 6.3, 6.4, 6.5). Using our maximalist computational approach, we resolve some features of the effective regulatory network, but we cannot completely distinguish between models given the data. We identify model sets, or models with shared regulatory features that generate acceptable fits to each dataset (see Figure 2 and Section 6.6). Similarly, we can rule out a host of possible regulatory network structures (see Figure 6). We find a highly interconnected regulatory network structure dominated by positive regulation that can explain the data across each of the three experimental conditions, suggesting a possible mechanism that underpins the mouth-form decision in Pristionchus pacificus. More work is needed to experimentally verify the proposed mechanism and, more broadly, refine our model sets. The results we find here could be used to guide experiments aimed at achieving a mechanistic understanding of regulation during developmental decisions. This study examines practical indistinguishability, a generalized form of sensitivity analysis across model structure given a data set, which is rarely considered in modeling studies but is increasingly important given the growing number of models being created. Our analysis allows us to probe the limits of what we can mechanistically infer, and is broadly applicable across scientific disciplines, including biology, chemistry, and physics.
@@ 3 Results
We tested representatives from our large family of models for structural identifiability using STRIKE-GOLDD, and found they were not structurally identifiable, due to the presence of a scaling symmetry [27]. Scaling symmetries are a type of algebraic transformation of parameters and variables of a model that leave the model invariant, meaning it produces the same dynamic behavior in the observables. Consider the following example model from our family of regulatory network models with the scaling symmetry highlighted in blue. From visual inspection, the scaling symmetry cancels from each equation of the model, meaning any value of $\xi$ produces the same mRNA dynamics.
@@
[EUD-1 protein], Unobs.: ${\color[rgb]{0,0,1}\xi}\dfrac{dE}{dt}={\color[rgb]{0,0,1}\xi}k_{t}M_{e}-{\color[rgb]{0,0,1}\xi}k_{-p}\hat{E}$, $E(0)=E_{o}$ (2)
@@
[NHR-40 protein], Unobs: ${\color[rgb]{0,0,1}\xi}\dfrac{dN}{dt}={\color[rgb]{0,0,1}\xi}k_{t}M_{n}-{\color[rgb]{0,0,1}\xi}k_{-p}N,$ $N(0)=N_{o}$ (4)
@@
[NHR-40 protein], Unobs: ${\color[rgb]{0,0,1}\xi}\dfrac{dN}{dt}={\color[rgb]{0,0,1}\xi}k_{t}M_{n}-{\color[rgb]{0,0,1}\xi}k_{-p}N,$ $N(0)=N_{o}$
@@
To eliminate the structural identifiability issue and break the scaling symmetry, we set $\xi=\dfrac{1}{k_{t}}$, which rescales the protein states and the half-max parameters by the translation rate. This is reasonable, as we know that the translation rate must be non-zero and has the benefit of reducing the dimension of the parameter space by one. We then checked a subset of the reparameterized models to ensure the resulting models were structurally identifiable using the package StructuralIdentifiability.jl [28]. For example, Eqs. (7)-(12) are a structurally identifiable version of Eqs. (1)-(6). We corrected all 13,824 models for this scaling symmetry ahead of parameter estimation (see Sections 6.3 and 6.4). Although we cannot verify the structural identifiability of all models due to computational limitations, removing at least this ubiquitous symmetry should improve the reliability of parameter estimation.
@@
Before fitting our family of models to the experimental data, we performed a series of synthetic tests using Eqs. (7)-(12) to understand the limits of inference in different data regimes. As Eqs. (7)-(12) are structurally identifiable, any parameter identifiability issues should be practical in nature, meaning they emerge from the data quality rather than from symmetries in the structure of the equations. To get a sense of how well we can recover the ground truth parameters, we sampled synthetic data from known parameters and varied the sampling frequency and noise level. We then refit the true model to the synthetic data sampled in different sampling and noise regimes. We also fit several “misspecified model structures,” or models with structural differences from the true model structure, to the synthetic data to understand the practical indistinguishability of Model 11574 from other models across different data regimes. We refer to practical distinguishability as the ability to uniquely infer the structure given the data. “Practically indistinguishable" was less examined in our search of the literature, perhaps due to a combination of its computational nature or the challenge of interpreting the results across multiple, competing model structures that emerge from structural uncertainty, though we did find some examples, such as [7].
@@
To determine how well we can recover ground truth parameters, we simulate synthetic data across a range of sampling rates corresponding to a few observations across the course of development to hourly sampling and a broad range of noise levels from Model 11574 (of 13,824) and attempt to fit the true model structure to the data (Figure 3, B top). We select this model because it generates a good fit to the observed data, which we discuss in subsequent sections. Model 11574 has the form
[eud-1 mRNA], Obs.: $\dfrac{dM_{e}}{dt}=\alpha_{1}+k_{t_{e_{1}}}\left(\dfrac{\hat{N}}{\hat{K_{1}}+\hat{N}}\right)\left(\dfrac{\hat{S}}{\hat{K_{4}}+\hat{S}}\right)-k_{-m}M_{e},$ $M_{e}(0)=M_{eo}$ (7)
[scaled EUD-1 protein], Unobs.: $\dfrac{d\hat{E}}{dt}=M_{e}-k_{-p}\hat{E}$, $\hat{E}(0)=\hat{E_{o}}$ (8)
[nhr-40 mRNA], Obs.: $\dfrac{dM_{n}}{dt}=\alpha_{2}+k_{t_{n_{1}}}\left(\dfrac{\hat{N}}{\hat{K_{2}}+\hat{N}}\right)-k_{-m}M_{n},$ $M_{n}(0)=M_{no}$ (9)
[scaled NHR-40 protein], Unobs.: $\dfrac{d\hat{N}}{dt}=M_{n}-k_{-p}\hat{N}$, $\hat{N}(0)=\hat{N_{o}}$ (10)
[sult-1 mRNA], Obs.: $\dfrac{dM_{s}}{dt}=\alpha_{3}+k_{t_{s_{1}}}\left(\dfrac{\hat{E}}{\hat{K_{3}}+\hat{E}}\right)^{2}-k_{-m}M_{s},$ $M_{s}(0)=M_{so}$ (11)
[scaled SULT-1 protein], Unobs.: $\dfrac{d\hat{S}}{dt}=M_{s}-k_{-p}\hat{S}$, $\hat{S}(0)=\hat{S_{o}}$, (12)
@@
To analyze the impact of experimental noise on the practical indistinguishability of this model from others, we needed to choose an error model. We observed that the noise was higher when the mean of the data was higher, potentially indicating the presence of signal-dependent noise, which is commonly observed in biology, for example, see [29]. Based on this, a multiplicative error model is the simplest reasonable model we can assume, as there was not enough data to empirically estimate the error distribution and there was no straightforward first-principles approach based on knowledge of the system. We assumed a data-centered Gaussian with standard deviation proportional to the mean and generated synthetic data for a wide range of sampling rates (row in Figure 3B) and noise levels (column in Figure 3B). More negative log-likelihood values (lighter colors) correspond to better fits (see Methods). Using this error model and the known sampling rate, we can approximate the experimental data regime (yellow box, Figure 3). When we fit the known model structure to synthetic data in the approximate experimental data regime, we get a reasonable fit to the data, though the exact parameter values are not recovered (Figure 3, B top, yellow box and C top). Based on these synthetic tests, we do not expect each individual model to have practically identifiable parameters.
@@
To understand if we can uniquely infer the structure of a given model from data sampled in the relevant experimental regime—or how practically distinguishable Model 11574 is from other model structures, we explore fitting misspecified models to the synthetic data. We select three alternative structures or misspecified models to fit to the synthetic data generated by Model 11574. We selected Models 7308, 353, and 1 of the 13,824 structures, as they vary in how closely related their structure is to Model 11574, the true model structure in this synthetic test (Figure 3 A). Model 7308 and Model 11574 share several regulatory features, whereas Models 353 and 1 have completely different regulatory structures from Model 11574 and vary in their complexity (see Figure 3A). Model 353 is much more complex than Model 1. In the relevant experimental data regime, we see that Model 7308 gives a reasonable fit to synthetic data generated from Model 11574 and has relatively tightly constrained uncertainty (Figure 3 C, second row), whereas Models 353 and 1 give a much poorer and more uncertain fit (Figure 3 C, third and fourth rows). In particular, the strongly mispecified models poorly describe the dynamics of nhr-40 mRNA over the entire course of development, as the uncertainty covers the entire possible range of the data. While we do not expect the parameters to be practically identifiable in the approximate experimental data regime, it is clear that some model structures are capable of generating a reasonable fit to the data, while others are not. Given the properties of the experimental data, we expect to be able to resolve some features of the network structure partially, but we do not expect this system to be practically distinguishable from all considered models.